OpenStudy (anonymous):

The sum of twice a number and 11 less than the number is the same as the difference btw -19 and the number. what is the number2:32 pm

OpenStudy (anonymous):

disregard the 2:32pm!

OpenStudy (anonymous):

look at mine after too

OpenStudy (anonymous):

If the number is x, then: \[2x+(x-11)=-19-x \implies 4x=-8 \implies x=-2\]

OpenStudy (anonymous):

how do you know all these! lol thanks

OpenStudy (anonymous):

2000 raffle tickets are sold for a dollar each. one prize is 1000 is to be awarded. find the fair price

OpenStudy (anonymous):

Haha.. I am a genius :P

OpenStudy (anonymous):

haha seriosuly keep helping me!!!

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

But you need to give me medals Gkouklan :P

OpenStudy (anonymous):

oh sorry im new to this site... how do i do that lol